OpenStudy (anonymous):

write the equation of the line which passes through (2,-3) and is perpendicular to y=4x+7 in standard form?

OpenStudy (saifoo.khan):

perp. line slope = -4

OpenStudy (anonymous):

i need steps

OpenStudy (saifoo.khan):

find the y-intercept now. -3 = 2(-4) + c c = -3 + 8 c = 5 therefore: y = 4x + 5

OpenStudy (anonymous):

m1 = 4, not -4/7.

OpenStudy (anonymous):

m1=4 for pependicular is m1xm2=-1 so m2=-1/4 y+3=-1/4(x-2) y=-1/4x-7/2

OpenStudy (anonymous):

sorry i meant this

OpenStudy (anonymous):

4y + x + 10 = 0